OpenStudy (anonymous):

solve the following equation 2 (h-8) -h = h-16

OpenStudy (john_es):

\[2h-2\cdot8-h=h-16\Rightarrow h-16=h-16 \Rightarrow 16=16\]It has infinite solutions.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

what about this one? 5+7x=11+7x

OpenStudy (john_es):

\[5-11=7x-7x\Rightarrow-6\neq0\] It has no solutions.